OpenStudy (anonymous):

Use the slope formula to find the missing coordinate. R(0, _), S(-4, 2) m = 5

pooja195 (pooja195):

Hi there! Welcome To OpenStudy So here's what we know \[ \huge~\rm~\bf\color{blue}{ \frac{2-?}{-4-0}=5} \] Any ideas on what "?" would be?

OpenStudy (anonymous):

18 22 -22 i have no idea

pooja195 (pooja195):

Are those the options?

OpenStudy (anonymous):

yes

pooja195 (pooja195):

Ok since you don't know we can plug in the options to figure out what statement would make the slope =5 2-18=?

OpenStudy (anonymous):

-16

pooja195 (pooja195):

-16/-4=?

OpenStudy (anonymous):

-12

pooja195 (pooja195):

We would be dividing and it would end up being 4

pooja195 (pooja195):

Which does not match up with the 5 as the slope s 18 is not the answer next 2-22=?

OpenStudy (anonymous):

-20

pooja195 (pooja195):

-20 divided by -4 =?

OpenStudy (anonymous):

5 s0 its 22

pooja195 (pooja195):

Yep :D

OpenStudy (anonymous):

thank you :)

pooja195 (pooja195):

you're welcome :)